OpenStudy (anonymous):

solve for x. 4(x - b) =x

OpenStudy (anonymous):

4x-4b=x -4b= -3x b=3/4x

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

You're welcome. Hopefully that was correct. Do you understand it?

OpenStudy (anonymous):

yes i qet it : D